# From Large to Super-Tiny: End-to-End Optimization for Cost-Efficient LLMs

### v2

@@ 3.3 Model Compression
To further compress model, we design two phases of pruning following NVIDIA [71], Llama3.2 [72], and Apple [73]: (1) Offline One-Shot Struct Pruning: Following Minitron[31], we explore two pruning strategies. For Depth Pruning, we compute layer importance scores and remove the $N$ least important layers based on the target pruning ratio. For Width Pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Section I.1; (2) Recovery training: The first step is General Instruction Fine-tuning via Knowledge Distillation. Adopting Minitron’s approach, we leverage logits from a larger general instruction teacher model for recovery training. Due to unavailability of the teacher’s original training data, we first perform model correction on our selected IFT dataset to stabilize distillation, followed by top-k logits distillation(Section 3.2). The second step is Domain-Specific Task Alignment. The pruned compact LM undergoes further aligning through knowledge distillation, producing a final domain-optimized LLM.
@@ 4.1 Experimental Setup
In our experiments, we utilize a large-scale dataset built from years of team experience and practice accumulated in the field. We reconstruct the intent classification and semantic slot framework within the real-world system into an API document specification, creating a function pool comprising 74 functional modules and 225 parameters. We sample 200k data points from a single business scenario as the training set and synthesize labels based on the LLM prototype outlined in Section 3.1. A combination of stratified and random sampling yields 11,922 samples from this scenario’s data as the in-domain test set. Additionally, we sample 7,867 data points using the same methodology from another business scenario, without some callable functions in the training set, as the out-of-domain test set. Compared to open-source datasets, our self-constructed dataset is more complex and closely aligned with real-world scenarios (see Table 1).
@@
During the training phases, base models are all from small instruct models (0.5B-7B) of the Qwen2.5 model family [11]. We conduct RFT on them and RL on 7B models. We perform pruning on 0.5B models and obtain Qwen2.5-0.4B-Width and Qwen2.5-0.4B-Depth, both under a 20% pruning ratio. Then, we transfer knowledge from 7B RFT-trained models to pruned and 0.5B models via KD. We also apply KD from 7B RL-trained models to 0.5B models, followed by another RL phase (denoted as KD*+RL). Finally, we use model quantization via W8A16/W4A16/W8A8 strategies on the KD-trained 0.5B model, detailed in Appendix E. The training process is cost-efficient, utilizing only 8 NVIDIA H20 GPUs. Computing resources are listed in Appendix F.
@@
We do evaluation by comparing our models with crowdsourced labelers and the BERT-based system consisting of two models based on BERT-large (330M)[75]. To evaluate the effectiveness of our method, we invite expert annotators alongside the LLM-as-a-judge-based metric AR (see Equation 6) for performance assessment. We also implement a retrieval module to select top 10 functions from the function pool, aiming to accelerate both training and inference phases, as detailed in Appendix D. In real-world scenarios, we deploy models on NVIDIA L20 GPU and perform effectiveness evaluation via TensorRT-LLM[74], with specific configuration details provided in Appendix H. During the comparison between the trained models and the BERT-based system, we adopt queries per second (QPS) as the primary metric for system throughput, response time (RT) for end-to-end latency, and concurrent clients (CC) to quantify high-concurrency performance.
@@ 4.2 Experimental Results
Table 2: Main results of the models trained on Qwen2.5(See Appendix J for FLOPs analysis).
Model / System AR GFLOPs
Crowdsourced Labelers 77.85% -
@@
Qwen2.5-0.4B-Width (KD) 94.17% 33.10
Qwen2.5-0.4B-Depth (KD) 94.20% 40.83
@@
Table 3: RFT vs RL on in-domain data and out-domain data.
@@
Qwen2.5-0.5B (KD*+RL) 95.55% 91.68%
Figure 4: Comparison over different models by professional annotators
@@
As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a win+tie rate of 95.59% over the BERT-based system and 92.98% over crowdsourced labelers. This demonstrates the superior performance of the LLM prototype. Through the evaluation via LLM-as-a-Judge, we find that the LLM Prototype surpasses the BERT-based System by 13.31% and crowdsourced annotations by 16.89%, as shown in Table 2. This conclusion is nearly consistent with that presented by professional annotators, demonstrating the effectiveness of the LLM judge.
@@
According to Table 3, while RFT performs well on in-domain data, it demonstrates worse results on out-of-domain data with a 3.61% AR decrease compared to the LLM prototype. Conversely, 7B RL-trained models achieve superior performance compared to other models on both in-domain and out-of-domain datasets, indicating better generalization of RL training methodology. Notably, even with models as small as 0.5B, employing a hybrid training methodology with KD+RL can yield superior in-domain results and an improvement of 1% in out-of-domain datasets compared to the RFT method. For more detailed exploration of the RL results, please refer to Appendix K.
Tiny LLMs as a Cost-Efficient Upgrade to BERT-based System
As shown in Table 2 and Table 4, quantization minimally impacts accuracy, with less than a 0.1% AR loss, while significantly enhancing inference performance, achieving up to 1.29x higher QPS and 0.76x lower RT in the optimal scenario. Pruning incurs a slightly higher 0.2% accuracy loss but consistently enhances performance by 20%. For end-to-end inference, with a CC of 64, the BERT-based system reaches 1110.88 QPS and 55.84 ms RT. Although the cost is roughly twice that of the BERT-based system, the AR improves by up to absolute 14%.
Table 4: Inference Performance under different model compression techniques.
@@
Figure 5: Normalized RT ratio based on FP16 models
@@ 5 Analysis
Our analysis identifies depth pruning as the optimal approach for domain-specific models. Depth pruning achieves 20% inference acceleration with lossless AR at same compression ratios (Table 2), stemming from the nature of depth pruning, removing entire transformer layers and their associated attention computations. Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 9), depth pruning presents optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix J). Regarding quantization, the W4A16 strategy achieves impressive performance in single-thread scenarios, while FP8 outperforms with lossless accuracy under high concurrency conditions as the superior paradigm for throughput deployments.
@@ 6 Conclusion
Based on our comprehensive study, several key insights emerge regarding the development and deployment of effective deep learning models in practical scenarios.
Tiny Models as a Practical Low-Cost Alternative
Our experiments indicate that end-to-end optimized tiny models can achieve performance levels comparable to their larger counterparts, offering resource efficiency without compromising accuracy. This provides an alternative pathway for applying LLMs in practical tasks under low-cost scenarios.
@@
Scalable Cost-Efficient Three-Stage Pipeline
The proposed pipeline-comprising prototyping, knowledge transfer, and model compression-presents a systematic framework designed to achieve high performance while maintaining low costs in both training and inference. We believe that this approach can be effectively adapted to other domains, thereby potentially advancing research in those areas and facilitating the widespread deployment of LLMs.
@@ 7 Limitation & Future Works
Our pipeline has presented optimal cost-performance results, and our super tiny models have exhibited remarkable capabilities, featuring low latency, cost efficiency, cloud independence, and robust privacy. Further compression may offer limited marginal benefits. Thus, we are more focused on enhancing such models’ capabilities and expand their broader applications such as mobile devices:
Hybrid Methodology for LLMs Capabilities Transfer
Based on our existing experiments, the mixer of reinforcement learning and knowledge distillation has effectively enhanced the generalization capabilities of super tiny models. This implies a potentially powerful hybrid methodology that allows super tiny models to nearly fully retain the capabilities of LLMs in domain-specific tasks. In future research, we will build upon our previous work to conduct deeper studies and theoretical analyses of this hybrid approach in transferring LLMs capabilities to such super tiny models.
On-device Deployment Optimization
Although we achieve 8-bit compression and online deployment, real-world edge devices face heterogeneous hardware constraints such as memory fragmentation and NPU compatibility. In the future, for on-device deployment of super tiny models, we will develop hardware-aware compression algorithms and adaptive runtime kernels to optimize latency-energy tradeoffs across diverse IoT ecosystems, and explore more resource-efficient deployment techniques for edge devices, such as sparse computation, edge-centric quantization, and federated learning.
Appendix A Dataset Analysis
Type-Token Rate (TTR), as a fundamental measure of lexical diversity, reflects the local variability of vocabulary, but it is sensitive to text length. The calculation formula is:
$TTR=\frac{N_{types}}{N_{tokens}}$ (7)
where $N_{tokens}$ denotes the total number of 2-gram patterns in the text, and $N_{types}$ represents the number of distinct 2-gram patterns.
Entropy, as an information-theoretic metric, quantifies the uncertainty within binary grammatical sequences based on their probability distribution. A higher entropy value indicates more complex local language patterns. The calculation formula is:
$Entropy=-\sum_{i}^{N_{types}}p_{i}\log_{2}(p_{i})$ (8)
where $p_{i}$ is the probability of occurrence for the $i^{th}$ 2-gram pattern.
Since both TTR and Entropy are sensitive to text length, to ensure a relatively fair comparison of dataset complexity, we randomly sample 700 entries from the test set of each dataset for calculation(as the ATIS dataset only contains 700 instances).
Appendix B Details for Reinforcement Learning
@@
Appendix C Prompt Compression
@@
Appendix D Retriever Setup
Initially, we select GTE-large [63], which is trained on the BERT-large [75], as our embedding model. Using this model, we generate embeddings of API descriptions produced by the LLMs. Then, we compare these embeddings with those of the API description from the document using cosine similarity. However, this model’s recall accuracy is not high. Subsequently, we switch to GTE-Qwen2-1.5B-instruct [63], a model trained on Qwen2-1.5B [76], which offers the advantage of fine-tuning for matching tasks through prompt modification. Next, utilizing this model with a customized prompt in Section D.1, we perform cosine similarity matching between the model-generated API descriptions and the JSON text of the entire API documentation, achieving a higher accuracy. During training, we optimize the API description model using rejection sampling, with implementation details provided in Section D.2. Finally, to achieve optimal inference acceleration, we train a classifier based on BERT-large [75] for top-1 recall using rejection sampling. The classifier takes Dialog, User Query, and necessary context as input and the function category as output. The functions corresponding to the top-k output logits are then considered as candidate functions for the subsequent model’s function call. This approach significantly accelerates inference due to the efficiency of BERT classifiers compared to autoregressive GPT models, while also enhancing function recall capabilities. However, it introduces scalability challenges since the recall model needs to be retrained with each update to the function pool. As illustrated in Table 5, this series of optimizations lead to a substantially improvement in API recall rates.
Table 5: Comparison of retrieval performance of models: GTE-large model, GTE-Qwen2-1.5B-instruct model, optimized API description model + GTE-Qwen2-1.5B-instruct model. Extensibility implies that the API pool can be dynamically modified.
@@ D.2 Retriever Optimization
Appendix E Training Details
Table 6: Training Parameters in different phases.
@@ Appendix E Training Details
Table 7: Different KD results on Qwen2.5-0.5B model.
@@
[…]
The main training parameters are detailed in Table 7. During the training phase, we employ the AdamW optimizer [77], setting the model’s maximum sequence length to $5120$. In the RL phase, we utilize the baseline variant of Reinforce++ [50] as the reinforcement learning algorithm, offering advantages in efficiency, conciseness, and training stability. Following [78], the K2 estimator is used instead of the original K1 in Reinforce++. In the AKL phase, the threshold parameter $\mu$ is reduced from 1 to 0.5. For pruning, we use the 0.5B-Instruct model as the base, implementing width/depth pruning strategies with a 20%/30% prune ratio, resulting in two models Qwen2.5-0.4B(0.35B)-Width and Qwen2.5-0.4B(0.35B)-Depth. These extremely compressed models are produced through a two-phase training process involving accuracy recovery training and knowledge distillation (KD). Comprehensive training details for the pruned models are available in Appendix I. For quantization, we primarily experiment with three strategies on the 0.5B model trained with AKL: W8A16, W4A16, and W8A8. Both W8A16 and W4A16 employ per-group symmetric int8 weight-only quantization with a group_size=128, while activations are in fp16. To improve accuracy, W4A16 use 512 training data samples combined with the GPTQ algorithm [33] for calibration. In contrast, W8A8 uses FP8 for both weights and activations. For the training frameworks, we utilize Megatron-LM [79] for SFT and KD, while OpenRLHF [80] is employed for RL. Data sampling during training is conducted using vLLM [81]. In the evaluation phase, we employ SGLang [82] to generate data.
Appendix F Computing resources
@@
Appendix G Detailed Analysis for Experimental Results for KD
@@
Appendix H Inference Details

### v3

@@ 3.3 Model Compression
To further compress model, we design two phases of pruning following NVIDIA [64], Llama3.2 [65], and Apple [66]: (1) Offline One-Shot Struct Pruning: Following Minitron[31], we explore two pruning strategies. For depth pruning, we compute layer importance scores and remove the $N$ least important layers based on the target pruning ratio. For width pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Appendix H; (2) Recovery training: The first step is General Instruction Fine-Tuning via Knowledge Distillation. Adopting Minitron’s approach, we leverage logits from a larger general instruction teacher model for recovery training. Due to unavailability of the teacher’s original training data, we first perform model correction on our selected IFT dataset to stabilize distillation, followed by top-k logits distillation(Section 3.2). The second step is Domain-Specific Task Alignment. The pruned compact LM undergoes further aligning through knowledge distillation, producing a final domain-optimized LLM.
@@ 4.1 Experimental Setup
In our experiments, we utilize a large-scale dataset built from years of team experience and practice accumulated in the field. We reconstruct the intent classification and semantic slot framework within the real-world system into an API document specification, creating a function pool comprising 74 functional modules with 225 arguments. We sample 200k data points from a single business scenario as the training set and synthesize labels based on the LLM prototype outlined in Section 3.1. A combination of stratified and random sampling yields 11,922 samples from this scenario’s data as the in-domain test set. Additionally, we sample 7,867 data points using the same methodology from another business scenario, without some callable functions in the training set, as the out-of-domain test set.
@@
During the training phases, base models are all from small instruct models (0.5B-7B) of the Qwen2.5 model family [11]. We conduct RFT on them and RL on 7B models. We perform pruning on 0.5B models and obtain Qwen2.5-0.4B-Width and Qwen2.5-0.4B-Depth, both under a 20% pruning ratio. Then, we transfer knowledge from 7B RFT-trained models to pruned and 0.5B models via KD. We also apply KD from 7B RL-trained models to 0.5B models, followed by another RL phase (denoted as KD*+RL). Finally, we use model quantization via W8A16/W4A16/W8A8 strategies on the KD-trained 0.5B model, detailed in Appendix D. The training process is cost-efficient, utilizing only 8 NVIDIA H20 GPUs. Computing resources are listed in Appendix E.
@@
We do evaluation by comparing our models with the BERT-based system consisting of two models based on BERT-large (0.33B)[68]. To evaluate the effectiveness of our method, we invite expert annotators alongside the LLM-as-a-Judge-based metric AR (see Equation 6) for performance assessment. We also implement a retrieval module to select top 10 functions from the function pool, aiming to accelerate both training and inference phases, as detailed in Appendix C. To further verify the effectiveness of our proposed approach, we train models on the xlam-function-calling-60k dataset generated by APIGen[69] and conduct validation using the Berkeley Function-Calling Leaderboard (BFCL)[70]. Details can be found in Appendix K. In real-world scenarios, we deploy models on NVIDIA L20 GPU and perform effectiveness evaluation via TensorRT-LLM[67], with specific configuration details provided in Appendix G. During the comparison between the trained models and the BERT-based system, we adopt queries per second (QPS) as the primary metric for system throughput, response time (RT) for end-to-end latency, and concurrent clients (CC) to quantify high-concurrency performance.
@@ 4.2 Experimental Results
Table 1: Main results of the models trained on Qwen2.5 (See Appendix I for FLOPs analysis).
Model / System AR FLOPs (G)
@@
Qwen2.5-0.4B-Width (KD) 94.17% 40.83
Qwen2.5-0.4B-Depth (KD) 94.20% 33.10
@@
Table 2: In-domain and out-of-domain performances of different knowledge transfer strategies.
@@
Qwen2.5-0.5B (KD*+RL) 95.99% 92.34%
Figure 4: Comparison of annotation results between the LLM prototype and the BERT-based system determined by professional annotators.
@@
As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a superior performance over the BERT-based system, with a win+tie rate of 95.59%. Through the evaluation via LLM-as-a-Judge, we find that the LLM Prototype surpasses the BERT-based system by 13.31%, as shown in Table 1. This conclusion is nearly consistent with that presented by professional annotators, demonstrating the effectiveness of the LLM judge.
@@
According to Table 2, while RFT performs well on in-domain data, it demonstrates worse results on out-of-domain data with a 3.61% AR decrease compared to the LLM prototype. Conversely, 7B RL-trained models achieve superior performance compared to other models on both in-domain and out-of-domain datasets, indicating better generalization of RL training methodology. Notably, even with models as small as 0.5B, employing a hybrid training methodology with KD+RL can yield superior in-domain results and an improvement of 1.7% in out-of-domain datasets compared to the RFT method. For more detailed exploration of the RL results, please refer to Appendix J.
Tiny LLMs as a Cost-Efficient Upgrade of BERT-based System
As shown in Table 1, Table 3 and Figure 5, quantization minimally impacts accuracy, with less than a 0.1% AR loss, while significantly enhancing inference performance, achieving up to 1.29x higher QPS and 0.76x lower RT in the optimal scenario. Pruning incurs a slightly higher 0.2% accuracy loss but consistently enhances performance by 20%. For end-to-end inference, with a CC of 64, the BERT-based system reaches 1110.88 QPS and 55.84 ms RT. Although the cost is roughly twice that of the BERT-based system, the AR improves by up to absolute 14%.
Table 3: Comparison of QPS metrics using different quantization and pruning strategies.
@@
Figure 5: Comparison of RT metrics using different quantization and pruning strategies.
@@ 5 Analysis
Our analysis identifies depth pruning as the optimal approach for domain-specific models. Depth pruning achieves 20% inference acceleration with lossless AR at same compression ratios (Table 1), stemming from the nature of depth pruning, removing entire transformer layers and their associated attention computations. Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 8), depth pruning presents optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix I). Regarding quantization, the W4A16 strategy achieves impressive performance in single-thread scenarios, while FP8 outperforms with lossless accuracy under high concurrency conditions as the superior paradigm for throughput deployments.
@@ 6 Conclusion
Based on our comprehensive study, several key insights emerge regarding the development and deployment of effective LLMs in practical scenarios.
A Scalable Practical Alternative for Cost-Efficient LLMs
Our proposed pipeline presents a systematic framework designed to achieve high performance while maintaining efficient costs in both development and deployment. Our end-to-end optimized tiny models also outperform their larger counterparts, offering resource efficiency without compromising accuracy. This provides an alternative pathway for applying LLMs in practical tasks under low-cost scenarios. We believe this approach can be effectively adapted to other domains, thereby potentially advancing research and facilitating the widespread deployment of LLMs.
LLM Capability Preservation via a Hybrid Strategy
Our hybrid strategy demonstrates strong effectiveness and practicality in preserving LLMs’ capabilities during knowledge transfer. It enhances generalization and high-quality CoT generation abilities via RL, safely retains these capabilities in smaller models via KD, and further improves the in-domain and out-of-domain performance of super-tiny models via another RL. This methodology judiciously and scientifically avoids irreversible loss of capabilities that might occur when directly deriving super-tiny models via RL. It progressively retains LLMs’ capabilities, transferring them step by step to super-tiny models.
@@ 7 Limitation & Future Works
Our pipeline has presented optimal cost-performance results, and our super-tiny models have exhibited remarkable capabilities, featuring low latency, cost efficiency, cloud independence, and robust privacy. Further compression may offer limited marginal benefits. Thus, we are more focused on enhancing such models’ capabilities and expand their broader applications.
Deeper Study on LLM Capability Transfer
Based on our experiments, the hybrid strategy effectively enhances super-tiny models’ generalization while almost entirely retaining LLMs’ capabilities in domain-specific tasks. In future research, we will conduct a deeper investigation and theoretical analysis of this potentially powerful methodology, building upon prior work to gain a better understanding of this hybrid strategy in transferring LLMs’ capabilities to super-tiny models.
Cross-domain Optimal Compression Exploration
It is evident that compressing models to 0.5B while retaining their capabilities across all domains is not feasible. The limitation boundary remains unclear. So, further research and exploration are required to better understand and define these limits.
Appendix A Details for Reinforcement Learning
@@
Appendix B Prompt Compression
@@
Appendix C Retriever Setup
Initially, we select GTE-large [61], which is trained on the BERT-large [68], as our embedding model. Using this model, we generate embeddings of API descriptions produced by the LLMs. Then, we compare these embeddings with those of the API description from the document using cosine similarity. However, this model’s recall accuracy is not high. Subsequently, we switch to GTE-Qwen2-1.5B-instruct [61], a model trained on Qwen2-1.5B [71], which offers the advantage of fine-tuning for matching tasks through prompt modification. Next, utilizing this model with a customized prompt in Section C.1, we perform cosine similarity matching between the model-generated API descriptions and the JSON text of the entire API documentation, achieving a higher accuracy. During training, we optimize the API description model using rejection sampling, with implementation details provided in Section C.2. Finally, to achieve optimal inference acceleration, we train a classifier based on BERT-large [68] for top-1 recall using rejection sampling. The classifier takes Dialog, User Query, and necessary context as input and the function category as output. The functions corresponding to the top-k output logits are then considered as candidate functions for the subsequent model’s function call. This approach significantly accelerates inference due to the efficiency of BERT classifiers compared to autoregressive GPT models, while also enhancing function recall capabilities. However, it introduces scalability challenges since the recall model needs to be retrained with each update to the function pool. As illustrated in Table 4, this series of optimizations lead to a substantially improvement in API recall rates.
Table 4: Comparison of retrieval performance of different models: GTE-large model, GTE-Qwen2-1.5B-instruct model, optimized API description model + GTE-Qwen2-1.5B-instruct model. Extensibility implies that the API pool can be dynamically modified.
@@ C.2 Retriever Optimization
Appendix D Training Details
Table 5: Training parameters of different knowledge transfer strategies.
@@ Appendix D Training Details
Table 6: Ablation of KD on Qwen2.5-0.5B models.
@@
Qwen2.5-0.5B (AKL[$\mu=0.5$]) 94.48%
Qwen2.5-0.5B (AKL[$\mu=0.7$]) 94.68%
Qwen2.5-0.5B (AKL[$\mu=0.9$]) 94.85%
The main training parameters are detailed in Table 6. During the training phase, we employ the AdamW optimizer [72] with the cosine learning rate schedule, setting the model’s maximum sequence length to $5120$. In the RL phase, we utilize the baseline variant of Reinforce++ [50] as the reinforcement learning algorithm, offering advantages in efficiency, conciseness, and training stability. Following [73], the K2 estimator is used instead of the original K1 in Reinforce++. In the AKL phase, the threshold parameter $\mu$ is reduced from 1 to 0.5. When training with data containing CoT, we set the number of training epochs to 2 and use the average of the FKL loss and the CE loss. For pruning, we use the 0.5B-Instruct model as the base, implementing width/depth pruning strategies with a 20%/30% prune ratio, resulting in two models Qwen2.5-0.4B(0.35B)-Width and Qwen2.5-0.4B(0.35B)-Depth. These extremely compressed models are produced through a two-phase training process involving accuracy recovery training and knowledge distillation (KD). Comprehensive training details for the pruned models are available in Appendix H. For quantization, we primarily experiment with three strategies on the 0.5B model trained with AKL: W8A16, W4A16, and W8A8. Both W8A16 and W4A16 employ per-group symmetric int8 weight-only quantization with a group_size=128, while activations are in fp16. To improve accuracy, W4A16 use 512 training data samples combined with the GPTQ algorithm [33] for calibration. In contrast, W8A8 uses FP8 for both weights and activations. For the training frameworks, we utilize Megatron-LM [74] for SFT and KD, while OpenRLHF [75] is employed for RL. Data sampling during training is conducted using vLLM [76]. In the evaluation phase, we employ SGLang [77] to generate data.
Appendix E Computing resources
@@
Appendix F Detailed Analysis for Experimental Results for KD
@@
Appendix G Inference Details